The probability that a student will pass chemistry class is 0.85. Three students are randomly selected. Find the probability that:
a) all 3 will pass the class
b) none of 3 will pass the class
c) at least 1 student will pass

Answer by ewatrrr(24785)

p(pass) = .85, n = 3
......
P(x = 3) = binompdf(3, .85, 3) 0r (.85)^3
P(x = 0) = binompdf(3, .85, 0) 0r (.15)^3 (Note: 1-.85 = .15)
P(x ≥ 1) = 1 - P(0) = 1 - (.15)^3